OpenStudy (anonymous):

cansomeone help. simplify (7)/(1-√3)

OpenStudy (anonymous):

@karan000 the answer choices are A. -(7-7√2)/(2 B. -(7+7√2)/(2) C. (4+6√3)/(3) D (2+7√3)/(2)

OpenStudy (mertsj):

Multiply numerator and denominator by the conjugate of the denominator.

OpenStudy (anonymous):

huh can you explain'

OpenStudy (mertsj):

\[\frac{7}{1-\sqrt{3}}\times \frac{1+\sqrt{3}}{1+\sqrt{3}}\]

OpenStudy (anonymous):

It appears all the answer choices are incorrect?

OpenStudy (anonymous):

so i multiply the top with the top and bottom with bottom or do i cross multiply? @karan000 idk im so lost

OpenStudy (mertsj):

You multiply two fractions by multiplying the numerators and multiplying the denominators.